Which equation represents a line perpendicular to the line whose equation is 2x+3y equals 12?
(1) 6y equals -4x+12 (3) 2y equals -3x+6
(2) 2y equals 3x+6 (4) 3y equals -2x+12
Please show me how you did it. Thank You

Answer by Cromlix(4381)

Sort the equation into y = mx + c
2x + 3y = 12 => 3y = -2x + 12 => y = -2/3x + 4.
This equation has a gradient = -2/3
For a line to be perpendicular to another,
their gradients multiply together to = -1
So if this equation's gradient = -2/3
then the other line's gradient = 3/2
-2/3 * 3/2 = -1
Check the gradients from (1) - (4)
1) -4/6 = -2/3
2) 3/2
3) -3/2
4) -2/3
Therefore (2) is the answer.
Hope this helps.
:-)